Solve 4x^3-28x^2+55x-21=0 using the decomposition method.

Decomposition method: (do not just factor)
Substitute x=a/3 to get in the form of y^3+py+q=0. Then find the discriminant value, etc...
Answer #1

Given polynomial equation is:

4x^3-28x^2+55x-21=0

4x^3-28x^2+48x+7x-21=0

4x^3-28x^2+48x+7(x-3)=0

4x^3-12x^2-16x^2+48x+7(x-3)=0

4x^3-12x^2-16x(x-3)+7(x-3)=0

4x^2(x-3)-16x(x-3)+7(x-3)=0

(4x^2-16x+7)(x-3)=0

(4x-14)(4x-2)(x-3)=0

2 (2x-7)(4x-2)(x-3)=0

(2x-7)(4x-2)(x-3)=0

2 (2x-7)(2x-1)(x-3)=0

(2x-7)(2x-1)(x-3)=0

Substitute x=a/3 In the polynomial equation, we get

4(\frac{a}{3})^3-28(\frac{a}{3})^2+55\frac{a}{3}-21=0

\frac{4a^3}{27}-\frac{28a^2}{9}+\frac{55a}{3}-21=0